OpenStudy (prince1342):

The function b(x) = 2x − 5 determines how many bags of dog food need to be purchased for an animal shelter, where x is the number of dogs at the shelter. The shelter manager uses m(b(x)) to find the amount of money to bring for the dog food purchase. The function m(x) = 3x + 1. Solve for how much money to bring when there are 10 dogs in the shelter.

OpenStudy (prince1342):

15 31 46 52

OpenStudy (olivia.f):

Plug 10 into the equation and you get 15. Use 15 to plug in next m(x) = 3x + 1 put x = 15 m(15) = 3(15) + 1 = ?